A cereal company states that the mean weight of a box of a certain cereal is 14 ounces. A consumer claims that the mean weight of cereal in its box is actually less than 14 oz. Test the consumer claim, if a sample of 50 such cereal boxes has mean 13.6 oz and SD 1.1 oz. Explain what your conclusion
Answer #1

The test hypothesis is

Now, the value of test static can be found out by following formula:

Since the sample size is n = 50, degrees of freedom on the t-test statistic are n-1 = 50-1 = 49

This implies that

Since, the t distribution is symmetric about zero, so

Since , we reject the null hypothesis in favor of the alternative hypothesis